OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4.

myininaya (myininaya):

f(x)=ax^3+bx^2+cx+d f(-2)=-4 f(-2)=-8a+4b-2c+d=-4 just find some coefficient that give -4 as the result for the above let a=1 b=1 c=2 d=1 so we have f(x)=x^3+x^2+2x+1

myininaya (myininaya):

wait it didn't work i made a mistake somewhere

myininaya (myininaya):

oh i see my mistake lol -8(1)+4(1)-2(2)+1 does not give -4 lol

myininaya (myininaya):

-8(1)+4(1)-2(2)+4=-4 right? so a=1 b=1 c=2 d=4

myininaya (myininaya):

f(x)=x^3+x^2+2x+4

myininaya (myininaya):

that works

myininaya (myininaya):

you could chose anynumbers satisfying -8a+4b-2c+d=-4 as long as a is not zero since it wants a polynomial of degree 3

OpenStudy (anonymous):

true

myininaya (myininaya):

a=1 b=1 c=0 d=0 f(x)=x^3+x^2 is another one

OpenStudy (anonymous):

do you know the quotient

myininaya (myininaya):

i mean f(x)=x^3+x^2+2x+4

myininaya (myininaya):

x^2-x+4

OpenStudy (anonymous):

okay okay hold sooo first i have too set up the synthetic division problem which is ??? :) im trying too do it step by step first

myininaya (myininaya):

what polynomial do you want to use?

OpenStudy (anonymous):

i can use the one you gave me but it has too be something like for example x^2-2x+4 by x-6

OpenStudy (anonymous):

myininaya sorry to bother you but can you have a look at my question when your done:) http://openstudy.com/groups/mathematics/updates/4dff1c2a0b8bbe4f12e764cb

myininaya (myininaya):

so you want to divide x^2-2x+4 by x-6? why are we doing this?

myininaya (myininaya):

okay and

myininaya (myininaya):

is this a separate problem?

myininaya (myininaya):

boo?

myininaya (myininaya):

6| 1 -2 4 | 6 24 ________________ 1 4 | 28 so x^2-2x+4 divided by x-6= x+4 +28/(x-6)

myininaya (myininaya):

this doesn't have anything to do with the initial problem you asked except x^2-2x+4 is the quotient if we chose f(x)=x^3+x^2+2x+4 where remainder is -4

OpenStudy (anonymous):

okay so we can choose the f(x)=x^3+x^2+2x+4 where the remainder is 4 but the same way the other example was solved thats how my teacher wants meh too solve this question that i am asking

myininaya (myininaya):

remainder is -4 and im sorry what is the question you are asking exactly?

OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4.

myininaya (myininaya):

we already this though

myininaya (myininaya):

a polynomial of degree 3 is f(x)=ax^3+bx^2+cx+d we are also given that it should satisy f(-2)=-4 f(-2)=a(-2)^3+b(-2)^2+c(-2)+d=-8a+4b-2c+d=-4 -8a+4b-2c+d=-4 you can choose and polynomial in the form ax^3+bx^2+cx+d such that when you do -8a+4b-2c+d you must get -4 you can choose any values for a, b, c, and d as long as it satisfies -8a+4b-2c+d=-4 and d does not equal zero since we want a poly of degree 3

OpenStudy (anonymous):

okay i get ittttttt